OpenStudy (anonymous):

Evaluate and write the answer in scientific notation:(4 x 10^5)÷(2.5 x 10^8) A. 1.6 x 105/8 B. 1.6 x 10-3 C. 1.6 x 103

Parth (parthkohli):

First here's a property. \(\Large \color{MidnightBlue}{\Rightarrow a^m \div a^n = a^{m - n} }\) Let's now divide 4 by 2.5. I'm getting 1.6. Lets now divide 10^8 and 10^3, that'll be 10^(8 - 3) = 10^5. Answer is 1.6 * 10^5.

Parth (parthkohli):

Or wait a second. It'd be 10^-3